# Evaluating crop rotations around the world using satellite imagery and causal machine learning

### v1

@@ A.4 Controlling for geography
To conduct these robustness checks, we first fit the same causal forests with the approach described in Section 4.4 and used the same seed for random number generation that was used for the analysis presented in the main text. We then used a modified version of the best_linear_projection function in the grf package [62] to fit each of the following three modifications of the model in Equation (3)
@@
A.5 Accounting for spatial correlations?
In this appendix, we discuss and explore whether it makes sense to account for spatial correlations in the analysis. The samples used in each analysis in the paper were relatively well spread out (Table A1), with median distance to the nearest sample being at least 1km in every causal forest analysis conducted. This suggests that the choice of whether or not one accounts for spatial correlations may not have much influence on the ultimate results and conclusions. As a sensitivity check, we test whether clustering the data based on 0.1° 0.1° grid cells and calculating standard errors accounting for such clustering leads to much larger confidence intervals. In this sensitivity analysis we find that the confidence intervals increase in size by at most a factor of 1.7.
@@ A.5 Accounting for spatial correlations?
To explore how accounting for spatial correlations impacts the final results, we considered each 0.1° 0.1° grid cell to be a unique cluster. We then repeated the causal forest analyses assessing precrop effects (Section 4.4), diversification effects (Section 4.6), and heterogeneity fo precrop effects with weather (Section 4.5), and set the “cluster” parameter causal_forest function based on the latitude and longitude of each sample and these 0.1° 0.1° grid cells. The average_treatment_effect function used to estimate treatment effects and the best_linear_projection used to estimate heterogeneity coefficients then automatically accounted for the clusters when calculating standard errors (as recommended by the grf package documentation [63], we used HC1 standard errors [71] for the latter for computational reasons and since they are similar to HC3 standard errors in large sample sizes).
Using clustered standard errors led to confidence intervals that were generally larger. In Figure A9, we show a histogram of the factors by which the confidence intervals grew when using clustered standard errors (relative to the confidence intervals in the main text). Overall, the confidence intervals did not grow substantially when accounting for spatial correlations using the aforementioned approach. Sometimes the confidence intervals shrunk by a small amount; however, we suspect this phenomenon was due to discrepancies with the random number generation in the causal forest analyses.
Appendix B Sampling Scheme
@@
Appendix C Removing outlier peak GCVI values
In China and France where Sentinel-2 was used, observations with peak GCVI less than or greater than were deemed as outliers and removed. In the US and Canada where Landsat was used, observations with peak GCVI less than or greater than were deemed as outliers and removed. These thresholds were determined by looking at histograms of the estimated peak GCVI values in each country. Figure A11 depicts histograms for each country of the peak GCVI values estimated via harmonic regression. In cases where points were removed due to being outliers with peak GCVI estimates above a threshold, a red dotted line depicts the threshold used.
@@ Appendix C Removing outlier peak GCVI values
Appendix D Irrigation Imputation Scheme
@@
Our results are likely not sensitive to our chosen scheme for imputing post-2017 irrigation classifications using pre-2018 classifications. This is because farmers in the US do not frequently switch irrigation status from year to year. The distribution of for samples in the US is plotted in Figure A12. It can be seen that is often either or , suggesting that switching irrigation status is indeed rare. Note that many of the cases where could be driven by measurement errors, as the irrigation data product had an overall accuracy of 90–97 depending on the region [65].
Appendix E Heterogeneity analysis in the presence of unmeasured confounders or mediators
In this appendix we show that even if the causal effect estimates provided in this paper are biased due to unmeasured variables, it is possible that our analysis of the heterogeneity of the rotation benefit with weather will not be biased. In particular, we present assumptions under which the causal effect estimates provided by a causal forest would be biased due to unmeasured confounders, but the estimates describing how those causal effects vary with weather would not be biased. An identical mathematical argument shows that if there are also unmeasured, manipulable mediators, then under similar assumptions, the causal forest gives biased estimates for the direct effects, but the estimates describing how those effects vary with weather would not be biased estimates of how the direct causal effects vary with weather (see Appendix E.4).
E.1 Notation and assumptions
@@
E.2 The desired causal estimands and the causal forest estimands
@@
E.3 Allowing for interactions between and in model for
@@
E.4 Setting where manipulable mediators are missing
@@
Appendix F A comparison of using a vegetation index rather than satellite-based crop yield estimates
@@
F.1 Description of SCYM estimates
@@
F.2 Imposable conditions that allow forgoing SCYM estimates
@@
F.3 A theoretical argument that SCYM and GCVI-based outcomes give the same results up to a scalar multiple
@@
F.4 Interpretation and implication of boxed result
@@
In spite of these caveats, the approach we present in this paper of using peak GCVI as an outcome variable rather than SCYM yield estimates can allow investigators to study the impacts of agronomic practices on yields without needing to run crop modeling simulations to calibrate the SCYM yield estimates. If an investigator has an idea of what GCVI time series-based summary statistic and weather covariates they would use to estimate crop yields for a particular crop type and region, they can ensure Assumptions (i) and (ii) are met and that forgoing fitting a SCYM model would only cost them an unknown scaling factor. This unknown scaling factor would be unnecessary if their goal is merely to assess directions and not magnitudes, or if they ultimately plan to calibrate their observational causal effect estimates with those from randomized experiments as in [12] or with subnational-level yield data as in Section 4.7.
F.5 An empirical comparison when using nonlinear yield maps.
@@
We compared the results of our causal forest analysis when using QDANN as the outcome variable versus when using normalized peak GCVI as the outcome variable (as is done in the main text of the paper). To create a head-to-head comparison, we ran the analyses using the same random seed for both outcome variable choices and when restricting our attention to the same exact samples (east of the 100° W line) such that both peak GCVI and QDANN estimates were available. We then rescaled the estimated effects that were based on peak GCVI to be in units of percent of the average QDANN yield using a county-level regression rescaling approach similar to that used in Section 4.7. In particular, for each of 7 rotations in the US, we fit a linear regression of county-level averages of QDANN yield estimates on county-level averages of normalized peak GCVI with sample weights according to the number of the rotation-specific overlap samples in each county (i.e., the sample weights for each county were based on the number of samples in each county such that the propensity score satisfied ). We then used the estimated slope coefficients from these regressions and the overall average QDANN yield of each crop type to convert the peak GCVI-based results to be in units of the percent of average estimated QDANN yield (of the outcome crop of interest). Finally, we also used the average QDANN yield estimates to convert results based on those yield estimates to also be in units of percent of the average QDANN yield (of the outcome crop of interest).
@@
In Figure A7, for each of 7 rotations in the US, we plot the estimated conditional average treatment effect (CATE) at each sample in the overlap region (i.e., we plot the values of for the samples such that ) when using peak GCVI as the outcome variable (x-axis) versus when using QDANN yield estimates as the outcome variable (y-axis). The correlations between the CATEs are quite high and moreover, lines through the origin fit the scatter plots quite well. This suggests that even though QDANN is a nonlinear approach to estimating yield, the theoretical conclusions from the previous subsections approximately hold in the sense that QDANN based results can be approximated by some unknown rescaling of peak GCVI-based results. In addition, the agreement between the best fit line through the origin (red) and the line through the origin with slope 1 (cyan) suggests that weighted regressions involving county-level averages (which were used in the main text and were used to rescale the x-axes in Figure A7) often give good approximations of this unknown rescaling factor.
@@
In Figure A8, for the 7 rotations where QDANN yield estimates were available, we plot the main outcomes studied in the text (estimated precrop effects, benefits of diversification, and rescaled heterogeneity coefficients with weather). The results are based on the previously mentioned US subsample and either use QDANN yield estimates (blue) or peak GCVI as the outcome variable (red, green). We consider two approaches for rescaling the GCVI-based results: one is to calibrate using the coefficients from a county-level regression of QDANN yield averages on peak GCVI averages (red) and the other is to use the slope of the best fit line through the origin from the scatter plots in Figure A7 (green). The results show that whether one uses QDANN yield estimates as an outcome variable or peak GCVI does not change the sign of the results and typically does not impact whether or not they are statistically significant. Results based on QDANN roughly differ from those based on peak GCVI by some scalar as evidenced by the agreement between the estimates plotted in green versus those plotted in blue. However, when county-level yield estimates are used for rescaling, the results based on QDANN can have lower magnitude.
@@
Appendix G Labels considered non-crop in each country
In the benefit of diversification analysis described in Section 4.6, we considered rotations of the form BAB to be control units and the diverse treated units to be all samples of the form DAB, where D was any croptype that was distinct from A and B. When D was not an actual crop, we did not include it in the sample of treated units. In particular, in cases where D had any of the following classifications given in Table A3, the sample was not included in the analysis.
@@ Appendix G Labels considered non-crop in each country
Appendix H Validating transferability of US-based GCVI to yield calibration
@@
To compare a US-based calibration with those from Canada and France, for each outcome crop and year , we let be the average yield in each CAR based on the Canadian survey data and denote the average yield in each Department based on the French survey data. Similarly, for each outcome crop and year , we let be the average normalized peak GCVI in each CAR in Canada and be the average normalized peak GCVI in each Department in France based on the sample described in Section 4.1. For each crop type we then fit the following weighted linear regressions
where the weights were proportional to the sample sizes in our satellite-based dataset. We let , , , and denote the slope and intercept coefficients learned from the above regressions and further let and be the weighted average of yield of crop type based on subnational-level yield data in Canada and France (using the same weights as used in the linear regressions). Finally we let and denote the calibration coefficients learned from the regression in Equation (4) in the main text and let be the weighted average of crop yield based on county-level yield data in the US described in Section 4.7.
@@
US-based calibration (non-normalized): Estimate , and with
respectively.
@@
US-based calibration (normalized): Estimate and with
respectively. By cancelation, estimate with .
@@
Local calibration: Estimate , and with
respectively.
We compared the accuracy of the three approaches for estimating , and by using a weighted RMSE, where the weights used for computing the RMSE were proportional to the number of samples in our satellite-based dataset. The weighted RMSEs for the 3 approaches are presented in Table A4.
@@
The results in Table A4 show that the first approach, which does not normalize by average US crop yields, leads to far worse performance than the second approach, which normalizes by average US crop yields. Meanwhile, we see that the third approach, which uses a locally fit calibration model, performs only slightly better than the second approach, which uses a normalized US-based calibration. Therefore, in the main text we use the second approach because of its superior performance to the non-normalized approach and because, unlike the local calibration approach which performs only slightly better, it can readily be used in countries such as China where we do not have access to subnational-level yield data at sufficiently high resolution.
We also remark that the weighted RMSEs in Table A4 for winter wheat yields in France are relatively high compared to those in the US. One explanation for this is that the French subnational-level yield data downloaded from [85] included spelt and soft winter wheat in the same category whereas our data sample from which we calculated the average normalized peak GCVI only included soft winter wheat. Thus, the weighted RMSE reported for winter wheat in France using the US-based calibration approach may be artificially high. Further, given that the subnational-level yield data in France does not distinguish between spelt and soft winter wheat, a local calibration with such data in France could be unreliable and less accurate than a US-based calibration.
Appendix I Checking fit of causal forests
We considered a few metrics to assess the goodness of fit of the fitted causal forest and propensity score models that were used in this study. The goodness of fit metrics are described below and presented in Table A7 for each causal forest that was used to study 2-year cropping sequences.
@@ Appendix I Checking fit of causal forests
The ECE values are reported in the final column of Table A7 and ranged from 0.005–0.027, suggesting the estimated propensity score function was well calibrated in all examples in the study.
@@
We computed these causal forest based estimates of on all the training samples and calculated the RMSE relative to the actual . The RMSE was then rescaled to be in units of % of the average yield for that particular outcome crop and country (using the approach in Section 4.7). The rescaled RMSEs are reported in the penultimate column of Table A7, and ranged from –.
@@
In the above linear regression, is the previously mentioned random forest-based estimate of , is the propensity score estimating , is the causal forest based estimate of defined in Equation (2), is the overlap weighted sample mean of (which gives the estimated ATO), and is mean error. In the above model, we would expect and to be near 1 if the causal forests are well calibrated. Indeed, in Table A7, we find that the estimates of and are near 1 (with the exception of the Winter WheatCorn rotation in the US and some of the rotations in Canada involving corn).
Overall, the results in Table A7 suggest that the fitted causal forests and propensity score models were well calibrated. The estimated values of and being near suggested that the causal forests were generally well calibrated and that the estimated rotation benefits were heterogeneous. The ECE metric being generally close to suggested that the propensity score models were well calibrated. The RMSE for the predicted yields ranging from – of the average yield do not suggest the causal forest model was poorly calibrated and merely suggested that the features and rotation status did not explain all of the observed variation in yield (or peak GCVI).
For the casual forests fit in Section 4.6, the RMSEs were a bit higher (ranging from –) and the ECE scores still suggested a well calibrated propensity score models (ranging from –). In some instances the causal forests used to assess diversification effects had estimates of and that were not near , likely due to small sample sizes, although because we did not study heterogeneity of diversification effects with weather, we suspect such cases of poor calibration were not problematic for our analysis as we merely considered mean effects of diversification.
Appendix J How our heterogeneity with weather findings relate to existing claims in the literature
@@
Appendix K Supplemental Tables and Figures
@@
| Multiple countries? | [12, 13, 14, 15, 16, 17, 18, 19, 20] | |
| More than 2 types of rotation studied? | [12, 13, 14] | [15, 16, 17, 18, 19, 20] |
| Stratifies based on precrop and outcome crop? | [19] | [12, 13, 14, 15, 16, 17, 18, 20] |
| Assesses benefit of diversification? | [12, 13, 14, 15, 16, 17, 18, 19, 20] | |
| Assesses heterogeneity with weather? | [13, 14, 15, 16, 18] | [12, 17, 19, 20] |
| Uses formal causal inference framework? | [13, 14, 15, 16, 17, 18] | [12, 19, 20] |
| Confidence intervals for estimated effects? | [12, 16, 17, 18, 20] | [13, 14, 15, 19] |
| Confidence intervals for weather interactions? | [12, 13, 14, 15, 16, 17, 18, 19] | [20] |

### v2

@@ C.4 Controlling for geography
To conduct these robustness checks, we first fit the same causal forests with the approach described in Section A.4 and used the same seed for random number generation that was used for the analysis presented in the main text. We then used a modified version of the best_linear_projection function in the grf package [38] to fit each of the following three modifications of the model in Equation (3)
@@
C.5 Accounting for spatial correlations?
In this appendix, we discuss and explore whether it makes sense to account for spatial correlations in the analysis. The samples used in each analysis in the paper were relatively well spread out (Table S2), with median distance to the nearest sample being at least 1km in every causal forest analysis conducted. This suggests that the choice of whether or not one accounts for spatial correlations may not have much influence on the ultimate results and conclusions. As a sensitivity check, we test whether clustering the data based on 0.1° 0.1° grid cells and calculating standard errors accounting for such clustering leads to much larger confidence intervals. In this sensitivity analysis we find that the confidence intervals increase in size by at most a factor of 1.7.
@@ C.5 Accounting for spatial correlations?
To explore how accounting for spatial correlations impacts the final results, we considered each 0.1° 0.1° grid cell to be a unique cluster. We then repeated the causal forest analyses assessing precrop effects (Section A.4), diversification effects (Section A.6), and heterogeneity fo precrop effects with weather (Section A.5), and set the “cluster” parameter causal_forest function based on the latitude and longitude of each sample and these 0.1° 0.1° grid cells. The average_treatment_effect function used to estimate treatment effects and the best_linear_projection used to estimate heterogeneity coefficients then automatically accounted for the clusters when calculating standard errors (as recommended by the grf package documentation [38], we used HC1 standard errors [73] for the latter for computational reasons and since they are similar to HC3 standard errors in large sample sizes).
Using clustered standard errors led to confidence intervals that were generally larger. In Figure S10, we show a histogram of the factors by which the confidence intervals grew when using clustered standard errors (relative to the confidence intervals in the main text). Overall, the confidence intervals did not grow substantially when accounting for spatial correlations using the aforementioned approach. Sometimes the confidence intervals shrunk by a small amount; however, we suspect this phenomenon was due to discrepancies with the random number generation in the causal forest analyses.
C.6 Grid cell-level difference-in-means precrop effect estimates
In this subsection, we consider an intuitive, heuristic approach for estimating the precrop effects for comparison with the more formal causal forest approach. We restrict our attention to four precrop and outcome crop combinations in the US (SoybeansCorn, CornSoybeans, SoybeansSpring Wheat, and SoybeansWinter Wheat).
For each precrop A and outcome crop B considered, we let for samples (pixel-year pairs) that had the crop sequence AB and let for samples (pixel-year pairs) that had the crop sequence BB, and removed all other samples from the analysis. We then partitioned the map of the US into 1°1° grid cells. Then for each 1°1° grid cell and each year that had at least 25 samples with (corresponding to an AB rotation) and at least 25 samples with (corresponding to a BB rotation), we took the average value of the normalized peak GCVI for the samples with and for the samples with . In particular, we let and be the average of the normalized peak GCVI in the rotated group and non-rotated group, respectively, across the samples in grid cell and year . A heuristic estimate of the precrop effect is , and these estimates were then converted to the percent of average yield scale using the approach described in Section A.7.
For four different rotations in the US, Figure S11 gives a histogram of these heuristic precrop effect estimates across all grid cells an years with at least 25 samples of both and . The histograms suggest large positive average precrop effects for the CornSoybeans and SoybeansSpring Wheat sequences in the US, a small positive average precrop effect for the SoybeansCorn sequence in the US, and substantial negative average precrop effect for the SoybeansWinter Wheat sequence in the US. These qualitative findings are consistent with the average estimated precrop effects estimated using the causal forest (see Table S9 (Column 6)).
Appendix D Sampling Scheme
@@
Appendix E Removing outlier peak GCVI values
In China and France where Sentinel-2 was used, observations with peak GCVI less than or greater than were deemed as outliers and removed. In the US and Canada where Landsat was used, observations with peak GCVI less than or greater than were deemed as outliers and removed. These thresholds were determined by looking at histograms of the estimated peak GCVI values in each country. Figure S13 depicts histograms for each country of the peak GCVI values estimated via harmonic regression. In cases where points were removed due to being outliers with peak GCVI estimates above a threshold, a red dotted line depicts the threshold used.
@@ Appendix E Removing outlier peak GCVI values
Appendix F Irrigation Imputation Scheme
@@
Our results are likely not sensitive to our chosen scheme for imputing post-2017 irrigation classifications using pre-2018 classifications. This is because farmers in the US do not frequently switch irrigation status from year to year. The distribution of for samples in the US is plotted in Figure S14. It can be seen that is often either or , suggesting that switching irrigation status is indeed rare. Note that many of the cases where could be driven by measurement errors, as the irrigation data product had an overall accuracy of 90–97 depending on the region [37].
Appendix G Heterogeneity analysis in the presence of unmeasured confounders or mediators
In this appendix we show that even if the causal effect estimates provided in this paper are biased due to unmeasured variables, it is possible that our analysis of the heterogeneity of the rotation benefit with weather will not be biased. In particular, we present assumptions under which the causal effect estimates provided by a causal forest would be biased due to unmeasured confounders, but the estimates describing how those causal effects vary with weather would not be biased. An identical mathematical argument shows that if there are also unmeasured, manipulable mediators, then under similar assumptions, the causal forest gives biased estimates for the direct effects, but the estimates describing how those effects vary with weather would not be biased estimates of how the direct causal effects vary with weather (see Appendix G.4).
G.1 Notation and assumptions
@@
G.2 The desired causal estimands and the causal forest estimands
@@
G.3 Allowing for interactions between and in model for
@@
G.4 Setting where manipulable mediators are missing
@@
Appendix H A comparison of using a vegetation index rather than satellite-based crop yield estimates
@@
H.1 Description of SCYM estimates
@@
H.2 Imposable conditions that allow forgoing SCYM estimates
@@
H.3 A theoretical argument that SCYM and GCVI-based outcomes give the same results up to a scalar multiple
@@
H.4 Interpretation and implication of boxed result
@@
In spite of these caveats, the approach we present in this paper of using peak GCVI as an outcome variable rather than SCYM yield estimates can allow investigators to study the impacts of agronomic practices on yields without needing to run crop modeling simulations to calibrate the SCYM yield estimates. If an investigator has an idea of what GCVI time series-based summary statistic and weather covariates they would use to estimate crop yields for a particular crop type and region, they can ensure Assumptions (i) and (ii) are met and that forgoing fitting a SCYM model would only cost them an unknown scaling factor. This unknown scaling factor would be unnecessary if their goal is merely to assess directions and not magnitudes, or if they ultimately plan to calibrate their observational causal effect estimates with those from randomized experiments as in [12] or with subnational-level yield data as in Section A.7.
H.5 An empirical comparison when using nonlinear yield maps.
@@
We compared the results of our causal forest analysis when using QDANN as the outcome variable versus when using normalized peak GCVI as the outcome variable (as is done in the main text of the paper). To create a head-to-head comparison, we ran the analyses using the same random seed for both outcome variable choices and when restricting our attention to the same exact samples (east of the 100° W line) such that both peak GCVI and QDANN estimates were available. We then rescaled the estimated effects that were based on peak GCVI to be in units of percent of the average QDANN yield using a county-level regression rescaling approach similar to that used in Section A.7. In particular, for each of 7 rotations in the US, we fit a linear regression of county-level averages of QDANN yield estimates on county-level averages of normalized peak GCVI with sample weights according to the number of the rotation-specific overlap samples in each county (i.e., the sample weights for each county were based on the number of samples in each county such that the propensity score satisfied ). We then used the estimated slope coefficients from these regressions and the overall average QDANN yield of each crop type to convert the peak GCVI-based results to be in units of the percent of average estimated QDANN yield (of the outcome crop of interest). Finally, we also used the average QDANN yield estimates to convert results based on those yield estimates to also be in units of percent of the average QDANN yield (of the outcome crop of interest).
@@
In Figure S8, for each of 7 rotations in the US, we plot the estimated conditional average treatment effect (CATE) at each sample in the overlap region (i.e., we plot the values of for the samples such that ) when using peak GCVI as the outcome variable (x-axis) versus when using QDANN yield estimates as the outcome variable (y-axis). The correlations between the CATEs are quite high and moreover, lines through the origin fit the scatter plots quite well. This suggests that even though QDANN is a nonlinear approach to estimating yield, the theoretical conclusions from the previous subsections approximately hold in the sense that QDANN based results can be approximated by some unknown rescaling of peak GCVI-based results. In addition, the agreement between the best fit line through the origin (red) and the line through the origin with slope 1 (cyan) suggests that weighted regressions involving county-level averages (which were used in the main text and were used to rescale the x-axes in Figure S8) often give good approximations of this unknown rescaling factor.
@@
In Figure S9, for the 7 rotations where QDANN yield estimates were available, we plot the main outcomes studied in the text (estimated precrop effects, benefits of diversification, and rescaled heterogeneity coefficients with weather). The results are based on the previously mentioned US subsample and either use QDANN yield estimates (blue) or peak GCVI as the outcome variable (red, green). We consider two approaches for rescaling the GCVI-based results: one is to calibrate using the coefficients from a county-level regression of QDANN yield averages on peak GCVI averages (red) and the other is to use the slope of the best fit line through the origin from the scatter plots in Figure S8 (green). The results show that whether one uses QDANN yield estimates as an outcome variable or peak GCVI does not change the sign of the results and typically does not impact whether or not they are statistically significant. Results based on QDANN roughly differ from those based on peak GCVI by some scalar as evidenced by the agreement between the estimates plotted in green versus those plotted in blue. However, when county-level yield estimates are used for rescaling, the results based on QDANN can have lower magnitude.
@@
Appendix I Labels considered non-crop in each country
In the benefit of diversification analysis described in Section A.6, we considered rotations of the form BAB to be control units and the diverse treated units to be all samples of the form DAB, where D was any croptype that was distinct from A and B. When D was not an actual crop, we did not include it in the sample of treated units. In particular, in cases where D had any of the following classifications given in Table S4, the sample was not included in the analysis.
@@ Appendix I Labels considered non-crop in each country
Appendix J Validating transferability of US-based GCVI to yield calibration
@@
To compare a US-based calibration with those from Canada and France, for each outcome crop and year , we let be the average yield in each CAR based on the Canadian survey data and denote the average yield in each Department based on the French survey data. Similarly, for each outcome crop and year , we let be the average normalized peak GCVI in each CAR in Canada and be the average normalized peak GCVI in each Department in France based on the sample described in Section A.1. For each crop type we then fit the following weighted linear regressions
where the weights were proportional to the sample sizes in our satellite-based dataset. We let , , , and denote the slope and intercept coefficients learned from the above regressions and further let and be the weighted average of yield of crop type based on subnational-level yield data in Canada and France (using the same weights as used in the linear regressions). Finally we let and denote the calibration coefficients learned from the regression in Equation (4) in the main text and let be the weighted average of crop yield based on county-level yield data in the US described in Section A.7.
@@
US-based calibration (non-normalized): Estimate , and with ^α^(j,US)+^λ^(j,US)¯V_c,t^(j,Ca), ^α^(j,US)+^λ^(j,US)¯V_c,t^(j,Fr), and ^α^(j,US)+^λ^(j,US)¯V_c,t^(j,US), respectively.
@@
US-based calibration (normalized): Estimate and with ( ^α(j,US)¯Y(j,US)+^λ(j,US)¯Y(j,US) ׯV_c,t^(j,Ca) ) ׯY^(j,Ca) and ( ^α(j,US)¯Y(j,US)+^λ(j,US)¯Y(j,US) ׯV_c,t^(j,Fr) ) ׯY^(j,Fr), respectively. By cancelation, estimate with .
@@
Local calibration: Estimate , and with ^α^(j,Ca)+^λ^(j,Ca)¯V_c,t^(j,Ca), ^α^(j,Fr)+^λ^(j,Fr)¯V_c,t^(j,Fr), and ^α^(j,US)+^λ^(j,US)¯V_c,t^(j,US), respectively.
We compared the accuracy of the three approaches for estimating , and by using a weighted RMSE, where the weights used for computing the RMSE were proportional to the number of samples in our satellite-based dataset. The weighted RMSEs for the 3 approaches are presented in Table S5.
@@
The results in Table S5 show that the first approach, which does not normalize by average US crop yields, leads to far worse performance than the second approach, which normalizes by average US crop yields. Meanwhile, we see that the third approach, which uses a locally fit calibration model, performs only slightly better than the second approach, which uses a normalized US-based calibration. Therefore, in the main text we use the second approach because of its superior performance to the non-normalized approach and because, unlike the local calibration approach which performs only slightly better, it can readily be used in countries such as China where we do not have access to subnational-level yield data at sufficiently high resolution.
We also remark that the weighted RMSEs in Table S5 for winter wheat yields in France are relatively high compared to those in the US. One explanation for this is that the French subnational-level yield data downloaded from [86] included spelt and soft winter wheat in the same category whereas our data sample from which we calculated the average normalized peak GCVI only included soft winter wheat. Thus, the weighted RMSE reported for winter wheat in France using the US-based calibration approach may be artificially high. Further, given that the subnational-level yield data in France does not distinguish between spelt and soft winter wheat, a local calibration with such data in France could be unreliable and less accurate than a US-based calibration.
Appendix K Checking fit of causal forests
We considered a few metrics to assess the goodness of fit of the fitted causal forest and propensity score models that were used in this study. The goodness of fit metrics are described below and presented in Table S8 for each causal forest that was used to study 2-year cropping sequences.
@@
The ECE values are reported in the final column of Table S8 and ranged from 0.005–0.027, suggesting the estimated propensity score function was well calibrated in all examples in the study.
@@
We computed these causal forest based estimates of on all the training samples and calculated the RMSE relative to the actual . The RMSE was then rescaled to be in units of % of the average yield for that particular outcome crop and country (using the approach in Section A.7). The rescaled RMSEs are reported in the penultimate column of Table S8, and ranged from –.
@@
In the above linear regression, is the previously mentioned random forest-based estimate of , is the propensity score estimating , is the causal forest based estimate of defined in Equation (2), is the overlap weighted sample mean of (which gives the estimated ATO), and is mean error. In the above model, we would expect and to be near 1 if the causal forests are well calibrated. Indeed, in Table S8, we find that the estimates of and are near 1 (with the exception of the Winter WheatCorn rotation in the US and some of the rotations in Canada involving corn).
Overall, the results in Table S8 suggest that the fitted causal forests and propensity score models were well calibrated. The estimated values of and being near suggested that the causal forests were generally well calibrated and that the estimated rotation benefits were heterogeneous. The ECE metric being generally close to suggested that the propensity score models were well calibrated. The RMSE for the predicted yields ranging from – of the average yield do not suggest the causal forest model was poorly calibrated and merely suggested that the features and rotation status did not explain all of the observed variation in yield (or peak GCVI).
For the casual forests fit in Section A.6, the RMSEs were a bit higher (ranging from –) and the ECE scores still suggested a well calibrated propensity score models (ranging from –). In some instances the causal forests used to assess diversification effects had estimates of and that were not near , likely due to small sample sizes, although because we did not study heterogeneity of diversification effects with weather, we suspect such cases of poor calibration were not problematic for our analysis as we merely considered mean effects of diversification.
Appendix L How our heterogeneity with weather findings relate to existing claims in the literature
@@
Supplemental Tables and Figures
@@
| Multiple countries? | [12, 13, 14, 15, 16, 17, 18, 19, 20, 21] | |
| More than 2 types of rotation studied? | [12, 13, 14, 20] | [15, 16, 17, 18, 19, 21] |
| Stratifies based on precrop and outcome crop? | [19] | [12, 13, 14, 15, 16, 17, 18, 20, 21] |
| Assesses benefit of diversification? | [12, 13, 14, 15, 16, 17, 18, 19, 20, 21] | |
| Assesses heterogeneity with weather? | [13, 14, 15, 16, 18, 21]† | [12, 17, 19, 20] |
| Uses formal causal inference framework? | [13, 14, 15, 16, 17, 18, 21] | [12, 19, 20] |
| Confidence intervals for estimated effects? | [12, 16, 17, 18, 20, 21] | [13, 14, 15, 19] |
| Confidence intervals for weather interactions? | [12, 13, 14, 15, 16, 17, 18, 19, 21] | [20] |
†[21] studied heterogeneity of rotation benefit with geographical location, discussing how it relates to weather conditions